OpenStudy (anonymous):

What is the solution to 4x²+24x+36=0

OpenStudy (anonymous):

4(x^2 + 6x + 8) = 4(x + 4)(x + 2), x = -2 , x = -4

OpenStudy (anonymous):

\[-B \pm \sqrt{b^2-4\times A \times C} \over 2\times A\] This is the easiest way to factor and you will always get the right answer

OpenStudy (anonymous):

thanx